OpenStudy (anonymous):

find lim [sin^2(-)/(-)(1+cos(-))] (-)->0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{\sin^2 x}{x(1+\cos x)}\]?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

agree.... 0...

OpenStudy (anonymous):

but how did u guys get 0

OpenStudy (anonymous):

the limit as x->0 for sin(x)/x is equal to 1 its a general rule of thumbs

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{\sin x}{x}=1\] \[\lim_{x \rightarrow 0}\frac{\sin^2(x)}{(x)(1+\cos(x))}\] \[\lim_{x \rightarrow 0}\frac{\sin(x)}{(x)}*\lim_{x \rightarrow 0}\frac{\sin(x)}{(1+\cos(x))}\] \[1*\lim_{x \rightarrow 0}\frac{\sin(x)}{(1+\cos(x))}\] sin(0)=0 cos(0)=1

OpenStudy (anonymous):

oh thanks. :)